How many grams of pure silver and how many grams of an alloy that is 65% silver should be. melted together to produce 56g of an alloy that is 80% silver?
Answer by richwmiller(17219)

x+.65(56-x)=56*.80
x+36.4-.65x=44.8
.35x+36.4=44.8
.35x=8.4
multiply by 100 to remove decimals
35x=840
x=24 grams of pure silver are added to 32 grams of 65% silver
check
1*24+.65(32)=56*.80